OpenStudy (anonymous):

Please help me with this double integral; I'll type it below... Please help me with this double integral; I'll type it below... @Mathematics

OpenStudy (anonymous):

\[\int\limits_{0}^{3}\int\limits_{0}^{1}ze ^{y}\sqrt{1-z ^{2}}dzdy\]

OpenStudy (anonymous):

Use integral substitution first to eliminate z, and then integrate

OpenStudy (anonymous):

6.33

OpenStudy (anonymous):

I only got \[-1/3 e ^{3}\] which is wrong!

OpenStudy (zarkon):

\[\frac{e^3-1}{3}\]

OpenStudy (anonymous):

no -1/3e^3 but 1/3e^3

OpenStudy (anonymous):

zarkon, how did you get that?

OpenStudy (anonymous):

Could you show what you got after the substitution?

OpenStudy (zarkon):

remember e^(0)=1 ...I'm betting you are missing that part

OpenStudy (anonymous):

Ha! Exactly. Thank you. I got so used to ignoring the 0s after integration.

OpenStudy (zarkon):

no problem